# Integrating Agent-Based and Compartmental Models for Infectious Disease Modeling: A Novel Hybrid Approach

### v1

@@ 6.2 Computational Efficiency
In Hybrid Model B, while the overall trend shows reduced CPU time with increasing weight of the compartmental model, the relationship is more complex. The initial increase in CPU time at low values of $w_{2}$ reflects the added complexity due to the attraction point and the increased interaction between the models. Despite this complexity, the hybrid model still demonstrates a net reduction in computational cost compared to a pure ABM, especially at higher values of $w_{2}$. This indicates that even with added spatial heterogeneity, the hybrid model can offer computational advantages, though the exact savings depend on the specific configuration and interaction dynamics.
@@ 6.3 Limitations and Future Directions
One limitation of the hybrid model, compared to a pure ABM, is the loss of information during each exchange between the two model components. Whenever population flow from M2 to M1 is modeled, new agents are created and their individual statuses are randomly distributed based on the population’s current state. Consequently, an agent that recently entered M2 as recovered may effectively leave as susceptible or infectious. Exploring alternative strategies for information exchange between the models and refining the coupling mechanism could potentially address this limitation and improve the accuracy of the hybrid model in capturing individual-level dynamics.
Overall, these findings provide valuable insights into the strengths and limitations of hybrid modeling approaches. They emphasize the importance of carefully considering the spatial configuration and interaction dynamics between model components to optimize both accuracy and computational efficiency. Future research could focus on developing more sophisticated coupling mechanisms and exploring the application of hybrid models to a broader range of infectious diseases and real-world scenarios.
@@ 7 Conclusion
In this paper, we addressed a significant research gap by introducing a novel spatial hybrid modeling approach that integrates a microscopic agent-based model (ABM) with a macroscopic compartmental ordinary differential equation (ODE) model. This topic has been scarcely explored in the existing literature on hybrid models. Our primary aim was to explore the potential of this hybrid model in reducing computational costs while maintaining accurate and detailed infection dynamics. We demonstrated how such coupling can be effectively achieved, highlighting the potential of spatial coupling in real-world epidemic modeling, advancing the understanding of hybrid systems, and setting a foundation for future research in this domain.
To evaluate the effectiveness of our approach, we utilized several performance metrics to detect coupling effects on the resulting infectious disease dynamics. Our experiment with two toy scenarios revealed that the outcomes of the hybrid model depend on the similarity between the integrated models. In both scenarios, the ODE component lacked the spatial resolution of the ABM-component. However, in the first scenario, where both model components described a homogeneous population, the hybrid model produced the same infectious disease dynamics as the reference model. In the second scenario, we introduced spatial heterogeneity within the ABM by enabling non-random agent movement and localized clustering around an attraction point. The corresponding hybrid model showed that disease dynamics were influenced by the location of the between-model interface relative to the attraction point.
To address this, we suggest a meta-population approach for the ODE-component, which could provide spatial resolution and heterogeneity without significantly increasing computational demands. We also noted that the extreme variations in agent density in our toy examples may not be as pronounced in real-world applications, potentially mitigating concerns about model discrepancies.
To focus our analysis on the coupling effects, we intentionally simplified the model components, maintaining the standard Susceptible, Infectious, and Recovered (S, I, R) categories in both model components. This simplification allowed us to evaluate the impact of spatial interactions without the added complexity of more elaborate disease states or interactions. However, increasing the number of categories could further highlight the computational advantages of the hybrid approach compared to a purely microscopic approach. Future research will explore more realistic scenarios to further validate and enhance our proposed hybrid model.
@@ 8 Implementation
The model is implemented in Python v3.9 [37] and the code is available on GitHub (http://github.com/iebos/hybrid_model1). The model includes three primary class objects: The Person class which creates an agent, the Model_ODE class which solves the ODEs and the Model_hybrid_realistic_continuous which creates the ABM and couples it to the ODE.
@@ 9 Acknowledgements
The work on the paper was funded by the German Ministry of research and education (BMBF) (project MODUS-Covid, grant number 031L0302C), and under Germany’s Excellence Strategy (MATH+ -The Berlin Mathematics Research Center, grant number EXC-2046/1, project no. 390685689, subproject EF45-4). The authors extend their gratitude to Kai Nagel and Kristina Maier for their valuable insights and comments on an early draft of this manuscript.
@@ Appendix A Appendix 1
Figure 8: Hybrid model C: Average results for 100 independent realizations per level of model 2 weight ($w_{2}$). Bars around points represent the standard deviation. (a) RMSE for the time series of infectious between the reference model and the hybrid model. (b) Peak time of the time series of infectious. (c) Peak height of the time series of infectious. (d) Fisher-Pearson coefficient of skewness of the time series of infectious. (e) FWHM of the time series of infectious. (f) CPU time in seconds.
Figures 8(a) - 8(f) show results for Hybrid Model C, where the attraction point is placed at the center of M1. For example, at the pure ABM at $w_{2}=0$, it is at the same location as in Hybrid Model B at $(x=4.5,y=4.5)$. However, at $w_{2}=0.5$, it is placed at $(x=2.25,y=4.5)$. Most metrics behave differently than in results for Hybrid Model A and B. Rather than a steady or irregular trend towards the value of the reference model, most metrics rapidly increase or decrease towards the direction of the reference model and exceed or precede the reference value around $w_{2}=0.8$, before regressing towards it again. Most notably, the RMSE increases from $w_{2}=0.2$ to $w_{2}=0.74$ before sharply dropping. Given that the location of the attraction point can affect the population density within M1 and also the overall population distribution between the model components, it is likely that the fitted parameters of the pure ABM are not appropriate under these different conditions. Further, while the attraction point remains within M1, the distance to the interface grows smaller with increasing $w_{2}$. Figure 9(a) shows that as a result, between-model interaction increases continuously rather than peaking around the middle as is the case in Hybrid Model B. Additionally, Figure 9(b) shows that the majority of the overall population is located within M1 for $w_{2}<0.85$. Therefore, for Hybrid Model C, the weight of M2 is not a sufficient indicator of the influence of M2 on overall population dynamics.
Figure 9: Changes in population metrics for Hybrid Model A, B, and C for different weights of M2. Results from 100 independent realizations. (a) Average percentage of population entering M2 per $t$. (b) Average percentage of overall population in M1 per $t$.

### v2

@@ 6.2 Computational Efficiency
In Hybrid Model B, while the overall trend shows reduced CPU time with an increasing weight of the compartmental model, the relationship is more complex. The initial increase in CPU time at low values of $w_{2}$ reflects the added complexity due to the attraction point and the increased interaction between sub-models. Despite this complexity, the hybrid model still demonstrates a net reduction in computational cost compared to a pure ABM, especially at higher values of $w_{2}$. This indicates that even with added spatial heterogeneity, the hybrid model can offer computational advantages, though the exact savings depend on the specific configuration and interaction dynamics.
@@ 6.3 Limitations and Future Directions
One limitation of the hybrid model, compared to a pure ABM, is the loss of information during each exchange between the two sub-models. Whenever population flows from M2 to M1, the hybrid model creates new agents and randomly distributes their individual statuses based on M2’s current state. Consequently, an agent that recently entered M2 as recovered may effectively leave as susceptible or infectious. Exploring alternative strategies for information exchange between sub-models and refining the coupling mechanism could potentially address this limitation and improve the accuracy of the hybrid model in capturing individual-level dynamics.
Overall, our findings provide valuable insights into the strengths and limitations of hybrid approaches. They emphasize the importance of carefully considering the spatial configuration and interaction dynamics between sub-models to optimize both accuracy and computational efficiency. However, real-world epidemic outbreaks often introduce incubation times, antibody decline, or age-dependent differences in susceptibility, infectiousness, and behavior, among others. While our focus on a simplistic susceptible-infected-recovered scenario allowed us to isolate and examine the effects of hybrid coupling, we recognize that it has inherent limitations in terms of real-world applicability. Nonetheless, the presented hybrid model could extend to epidemic models with more realistic compartment structures and corresponding additions in the ABM. As this requires that disease-specific parameters and interaction patterns are integrated with rigor, modelers need to carefully consider the trade-off between the development time of the hybrid model and its benefits in terms of cost savings and accuracy over a single model paradigm approach.
Further, our toy model does not take full advantage of computationally more efficient ABM frameworks, making generalizations in terms of computational cost savings difficult. Given the experimental nature of our study, we limited our experiments to a small scale, focusing on the coupling mechanism and its implications for disease dynamics. In this case, modeling a larger population with our toy model would likely favor the hybrid approach even more, as the computational demands of the ODE component remain constant regardless of population size. We aimed to balance this potential discrepancy by modeling a relatively small population.
Future research could develop more sophisticated coupling mechanisms and explore scaling up this hybrid model using advanced architectures, such as cloud computing with MPI [42], high-performance platforms [33], or database routing techniques [32], to apply the model across a broader range of infectious diseases and real-world scenarios. These approaches could address the computational demands of large agent populations, presenting a promising direction for further exploration.
Another area of exploration involves the effect of population sizes and model configurations. Stochastic effects in small populations can create discrepancies between deterministic compartmental models and stochastic ABMs [43]. For infectious disease models, Rahmandad and Sterman [40] note that differences in model results are largely independent of population sizes ranging from 50 to 800, particularly when $R_{0}$ is high. Conversely, Hunter et al. [22] observed a stronger correlation between population size and outcomes in compartmental models compared to ABMs. Notably, differences in time discretization between these models are frequently overlooked, contributing to variations in results [21]. In our experiments, population sizes within model components varied based on the assigned component weights. Future research should examine the effects of time discretization and parameter variation in hybrid models while isolating the influence of population size differences on model consistency.
@@ 7 Conclusion
In this paper, we addressed a significant research gap by introducing a novel hybrid modeling approach that spatially integrates a microscopic agent-based model (ABM) with a macroscopic compartmental ordinary differential equation (ODE) model. This topic has been scarcely explored in the limited literature on hybrid models, and most studies lack clear descriptions of model architecture and systematic investigations on result consistency. Our primary aim was to explore the feasibility of this hybrid model in reducing computational costs while maintaining accurate and detailed infection dynamics. We demonstrated how such coupling can be effectively achieved, highlighting the potential of spatial coupling in real-world epidemic modeling, advancing the understanding of hybrid systems, and setting a foundation for future research in this domain.
To evaluate the effectiveness of our approach, we utilized several performance metrics to detect coupling effects on the resulting infectious disease dynamics. Our experiment with two toy scenarios revealed that the outcomes of the hybrid model depend on the similarity between the integrated models. In both scenarios, the ODE component lacked the spatial resolution of the ABM component. However, in the first scenario, where both sub-models describe a homogeneous population, the hybrid model produced the same infectious disease dynamics as the reference model. In the second scenario, we introduced spatial heterogeneity within the ABM by enabling landscape-driven agent movement and localized clustering around an attraction point. In the corresponding hybrid model, disease dynamics were influenced by the location of the between-model interface relative to the attraction point.
To address this, we suggest a meta-population approach for the ODE-component, which could provide spatial resolution and heterogeneity without significantly increasing computational demands. We also acknowledge that the extreme variations in agent density in our toy examples may not be as pronounced in real-world applications, potentially mitigating concerns about model discrepancies.
To focus our analysis on the coupling effects, we intentionally simplified the model components, maintaining the standard susceptible, infectious, and recovered categories in both sub-models. This allowed us to evaluate the impact of spatial integration without the added complexity of more elaborate disease states or interactions. However, increasing the number of categories could further highlight the computational advantages of the hybrid approach compared to a purely microscopic approach. Future research will explore more realistic scenarios and model configurations to further validate and enhance our proposed hybrid model.
@@ 8 Implementation
The model is implemented in Python v3.9 [44]. The project repository is available on GitHub (git.zib.de/ibostanc/hybrid_abm_ode), including the model code and jupyter notebooks for a tutorial and the presented experiments. The model involves three primary class objects: persons creates an agent, model_ode solves the ODEs, and hybrid_model creates the ABM and couples it to the ODE.
@@ 9 Acknowledgements
This work was funded by the German Ministry of research and education (BMBF) (project MODUS-Covid, grant number 031L0302C), and under Germany’s Excellence Strategy (MATH+ -The Berlin Mathematics Research Center, grant number EXC-2046/1, project no. 390685689, subproject EF45-4). The authors extend their gratitude to Kai Nagel and Kristina Maier for their valuable insights and comments on an early draft of this manuscript.
@@ Appendix A Appendix 1
Figure 8: Hybrid Model C: Average results for 100 independent realizations per level of M2 weight ($w_{2}$). Bars around points represent the standard deviation. (a) RMSE for the infectious time series between the reference model and the hybrid model. (b) Peak time of the infectious time series. (c) Peak height of the infectious time series. (d) Fisher-Pearson coefficient of skewness of the infectious time series. (e) FWHM of the infectious time series. (f) CPU time in seconds.
Figures 8(a) - 8(f) show results for Hybrid Model C with the attraction point centered in M1. For example, at $w_{2}=0$ (pure ABM), the location matches Hybrid Model B at $(x=4.5,y=4.5)$, while at $w_{2}=0.5$, it is placed at $(x=2.25,y=4.5)$. Most results differ from Hybrid Model A and B. Instead of a steady or irregular progression towards the reference model, most metrics initially converge towards the reference model but exceed/precede the reference value around $w_{2}=0.8$ before converging again. Notably, the RMSE increases from $w_{2}=0.2$ to $w_{2}=0.74$ before sharply dropping. Given the relationship between the attraction point location and the population density within M1 (and the distribution of the total population between components), the fitted parameters of the pure ABM are likely not suitable across different conditions. Further, while the attraction point remains within M1, the distance to the interface shrinks with increasing $w_{2}$. As depicted in Figure 9(a), this drives a steady rise in between-model interaction, contrasting with Hybrid Model B’s peak around $w_{2}=0.5$. Additionally, Figure 9(b) shows that the majority of the total population is located within M1 when $w_{2}<0.85$. Thus, for Hybrid Model C, $w_{2}$ alone does not a adequately reflect the influence of M2 on overall population dynamics.
Figure 9: Changes in population metrics for Hybrid Model A, B, and C for different M2 weights ($w_{2}$). Results from 100 independent realizations. (a) Average percentage of population entering M2 per $t$. (b) Average percentage of total population in M1 per $t$.